OpenStudy (anonymous):

Determine the axis of symmetry and the vertex of the given function. y = 2x2 − 12x + 21

OpenStudy (welshfella):

Axis of sYmmetry is x =-b/2a

OpenStudy (welshfella):

A b and c A

OpenStudy (anonymous):

who do you find the vertex?

OpenStudy (welshfella):

Where a b c are the values in equation ax2 +bx+c

OpenStudy (welshfella):

Plug the lie of x u just found into the function and compute y Vertex is the point (x,y)

OpenStudy (welshfella):

* value of x

OpenStudy (welshfella):

-b/2a = -(-12)/(2*2)=?